The random variable X represents the number of taxi cabs a doctor meets on his way to the hospital. The probability distribution of X is given in the following distribution table.



X P(x)


0 0.01


1 0.15


2 0.20


3 0.25


4 0.30


5 0.09



Find the expected value and variance of X.

1
Expert's answer
2022-03-03T09:37:24-0500
"mean=E(X)=\sum_ix_ip(x_i)"

"=0(0.01)+1(0.15)+2(0.20)+3(0.25)"

"+4(0.30)+5(0.09)=2.95"

"E(X^2)=0^2(0.01)+1^2(0.15)+2^2(0.20)"

"+3^2(0.25)+4^2(0.30)+5^2(0.09)=10.25"

"Var(X)=\sigma^2=E(X^2)-(E(X))^2"

"=10.25-(2.95)^2=1.5475"